Find the equation,show all work.
1. slope= -5/7 y int = -11
2. m= -1/2, (5,2)
3. m= 2/3, (6,5)

Answer by checkley77(12844)

Y=mX+b where m=slope & b=the Y intercept.
1. slope= -5/7 y int = -11
Y=-5X-11 ans.
2. m= -1/2, (5,2)
2=5(-1/2)+b
2=-5/2+b
b=2+5/2
b=2+2.5=4.5 ans. for the Y intercept.
Y=-X/2+4.5 ans.
3. m= 2/3, (6,5)
5=2/3*6+b
5=4+b
b=5-4=1 the Y intercept.
Y=2X/3+1 ans.